A dental office is in the process of changing their cost system. Currently use a single direct cost pool (professional labour) and a single indirect cost pool (staff support). Direct categories in new system include:
1. Professional partner labour. Avg compensation of 2 partners is $100,000 each, and each has 2000 hours of budgeted time
2. Dental assistant labour. Avg compensation of 4 assistants is $22,500 each, and each has 2000 hours budgeted time
3. Office staff. Avg compensation of 2 staff members is $15,000 each, and each has 2000 hours budgeted time.

Indirect category in new system includes professional liability insurance. Budgeted indirect amount is $200,000, and allocation base is budgeted professional hours. Dentist and dental assistant are considered professional labour hours

What is budgeted direct cost rate per hour for professional partner labour?

What is the budgeted indirect cost allocation rate per unit of the allocation base for the professional liability insurance?
1
Expert's answer
2016-04-23T06:43:05-0400
Budgeted direct cost rate per hour for professional partner labour is:
Direct cost = (100,000*2 + 22,500*4 + 15,000*2)/2,000 = 320,000/2,000 = $160 per hour.
The budgeted indirect cost allocation rate per unit of the allocation base for the professional liability insurance is:
Indirect cost = 200,000/2,000 = $100 per hour.